OpenStudy (anonymous):

A video camera is stationed 50 feet from a road that is part of a parade route to film the passing parade. Assume the parade floats are instructed to travel the route at a speed of 15 mi/hr. a) At what rate must the camera rotate to follor the parade floats when they are 25 feet from the point P on the road closest ot the camera? at the instant they pass point p? b) At what rate must the camera's focus adjust to film the floats clearly (that is, at what rate does the distance from the float to the camera change) when the floats are 200 feet from point P? when they are 50 feet from point P?

OpenStudy (anonymous):

I'm just not sure how to set it up.

OpenStudy (anonymous):

I'm awful with word problems.

OpenStudy (anonymous):

Okay, so one leg of the triangle changes at 15 mi/hr. The other leg is constantly 50. The first question is asking the rate of change of the angle.

OpenStudy (anonymous):

So write an equation relating the angle to the legs you know. The leg changing (we'll call it x) is opposite of the angle you care about, we'll call it angle a. Tan(a) = x/50

OpenStudy (anonymous):

oh so maybe use trig like sintheta=opp/hypp or something like that?

OpenStudy (llort):

is the road straight? how far down the road do you start filming?

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

That doesn't matter. Don't worry.

OpenStudy (anonymous):

But tan(a) = x/50, so solve for a. a = tan-1(x/50)

OpenStudy (anonymous):

Take the derivative of a with respect to x.

OpenStudy (anonymous):

(tan-1) is inverse tangent.

OpenStudy (anonymous):

ok. I just don't understand the triangle in that problem between the camera, road, and?

OpenStudy (anonymous):

I'll draw a picture. Just a moment.

OpenStudy (llort):

I am assuming that the line point p and the camera makes are perpendicular to the road, creating a right angle. so it would be tan^-1 (25/50) the other one would be tan^-1(200/50)-tan^-1(50/50)

OpenStudy (anonymous):

ok so da/dt=1/sqrt(1+(x/50)^2)

OpenStudy (anonymous):

times 1/50

OpenStudy (anonymous):

oh ok. I get it.

OpenStudy (anonymous):

Hope that picture clarifies.

OpenStudy (anonymous):

times dx/dt...forgot that part.

OpenStudy (anonymous):

Not quite right, I don't think. I have that derivative of arctan(x) is 1/(1+x^2). So derivative of arctan(x/50) would be (1/(1+x^2/50^2))*(1/50)

OpenStudy (anonymous):

no dx/dt at the end?

OpenStudy (anonymous):

The (1/50) is because of chain rule.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Yes, dx/dt also. That just gave you da/dx. You have to look at how quickly x is changing too.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not sure how to do that?

OpenStudy (anonymous):

I can't figure out a second equation that doesn't have too many variables.

OpenStudy (anonymous):

Okay, so we have da/dx = 1/(50 + x^2/50)

OpenStudy (anonymous):

So da/dt = da/dx * dx/dt, but dx/dt is given. It's 15 mi/hr.

OpenStudy (dumbcow):

i think it would be easier to convert that speed to ft per sec 15 mph = 22 ft/sec

OpenStudy (anonymous):

oh ok, I meant x :P not dx/dt.

OpenStudy (anonymous):

Good call, cow.

OpenStudy (anonymous):

but I'm guessing for a its 25.

OpenStudy (anonymous):

Yes. For a, it asks you to give da/dt when x = 25 and when x=0.

OpenStudy (anonymous):

Do you see how to use the equation we derived to solve those two things?

OpenStudy (dumbcow):

i get \[\frac{da}{dt} = \frac{(50)(22)}{x^{2} + 50^{2}}\]

OpenStudy (dumbcow):

also you need to know if they want it in radians per sec or degrees per sec

OpenStudy (anonymous):

I got da/dt=(1/(50sqrt(1+(x/50)^2))(dx/dt)

OpenStudy (anonymous):

I know, it looks super ugly.

OpenStudy (anonymous):

isn't it already in radians?

OpenStudy (anonymous):

Really not sure where your sqrt is coming from.

OpenStudy (dumbcow):

hmm yeah it should be, nevermind then

OpenStudy (anonymous):

derivative of arctan(x) is 1/sqrt(1+x^2)

OpenStudy (dumbcow):

no sqrt, just 1/(1+x^2)

OpenStudy (anonymous):

No it isn't?

OpenStudy (anonymous):

woops?

OpenStudy (anonymous):

Haha yes. Woops. You just used the wrong equation though. Everything else you did right.

OpenStudy (anonymous):

lol ok. so its that but w/o the sqrt?

OpenStudy (anonymous):

So your result should be da/dt = 1/(50 + x^2/50) * dx/dt

OpenStudy (anonymous):

Yes. That turns out to be what you put except without the sqrt.

OpenStudy (anonymous):

oh awesome!

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Need help on b?

OpenStudy (anonymous):

(My pleasure) =)

OpenStudy (anonymous):

yes please.

OpenStudy (anonymous):

Okay, so back to the picture, call that blue line h, for hypotenuse. We want to know how quickly it is changing with respect to time for the given x values.

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so opposite is 200 ft.

OpenStudy (anonymous):

So write an equation relating h to x.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

sintheta=200/h

OpenStudy (anonymous):

so theta=(sin-1)(200/h)

OpenStudy (anonymous):

Except that theta is going to be changing along with x, which would give us 2 variables. Can we write it more simply? Write an equality between h and x so that x is the only variable?

OpenStudy (anonymous):

If I give you x, how could you find h?

OpenStudy (anonymous):

oh duh h^2=50^2+200^2

OpenStudy (dumbcow):

thats correct but you dont want to use 200 yet, that is an x value you will plug in at the end you are looking for dh/dt you solved for da/dt dh/dt = da/dt * dh/da get h in terms of a first

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

Yes, except keep it as x for now.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

So h = sqrt(50^2 + x^2). Now find dh/dx.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

ok. dh/dt=(2x/2h)(dx/dt)

OpenStudy (dumbcow):

oh ok im wrong then, you only need dh/dx not dh/dt

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

Um... no you do need dh/dt. But first you should find dh/dx.

OpenStudy (anonymous):

so dh/dx=2x/2h

OpenStudy (anonymous):

But I'm not sure how you got that...

OpenStudy (anonymous):

h^2=50^2+x^2

OpenStudy (anonymous):

Derivative with respect to x of sqrt(50^2 +x^2)

OpenStudy (anonymous):

2h(dh/dx)=0+2x

OpenStudy (anonymous):

dh/dx=2x/2h

OpenStudy (anonymous):

First solve for h, so sqrt both sides.

OpenStudy (anonymous):

did I do something wrong? I was trying to not use the chain rule lol.

OpenStudy (anonymous):

Yeah, that's not quite right. It doesn't work that way.

OpenStudy (anonymous):

Okay, so you got to h^2 = 50^2 + x^2 and then tried to derive, but d/dx of h^2 is not 2h.

OpenStudy (anonymous):

d/dh of h^2 is 2h, but not d/dx.

OpenStudy (anonymous):

k. so dh/dx=1/2(2500+x^2)^-1/2(2x)

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

Yes. And (2500 + x^2)^-1/2 is 1/sqrt(2500 + x^2)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

So simplifying gives you 2x/(2*sqrt(2500+x^2))

OpenStudy (anonymous):

can I plug in x now?